OpenStudy (anonymous):

The sale of sweaters in a town changes with the temperature. If a statistical model predicts that daily sales (in dollars) are determined by the function s = 900 − 100t, where t is the temperature (in degrees Celsius), what is the temperature if sweater sales are at $1,200? –3°C 0°C –4°C 3°C

OpenStudy (saifoo.khan):

s = 900 - 100t. s = sales = 1200. Substitute in the function.

OpenStudy (anonymous):

1200=900-100t

OpenStudy (anonymous):

and I gets the -3 cause Thats what I gets

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

and it is -3

OpenStudy (saifoo.khan):

Yes.